OpenStudy (anonymous):

Pre-Cal! Can someone please help me convert 63°32’47” into a decimal degree form??

OpenStudy (mrnood):

there are 60 minutes in a degree and 60 seconds in a minute - so there are 3600 seconds in a degree You need to break up th ecalulation 47" = 47/3600 deg 32' = 32/60 deg Calculate those as decimals and add them up with the whole degrees

OpenStudy (anonymous):

would it be .54603?

OpenStudy (anonymous):

I'm sorry, .5463*

OpenStudy (mrnood):

just use the method above - I haven't done the calculation You need to add the whole degrees

OpenStudy (anonymous):

what do you mean by adding thee whole degrees?

myininaya (myininaya):

number in front of decimal

OpenStudy (anonymous):

there was no number in front though?

myininaya (myininaya):

yes there is unless I'm not suppose to see the 63 part.

OpenStudy (anonymous):

so it would be 63.5463?

myininaya (myininaya):

well i would go with 63.546 I think you are suppose to normally round to three decimals with these problems

myininaya (myininaya):

three decimal places *

OpenStudy (anonymous):

thank you so much! :)